OpenStudy (anonymous):

Let f(x) be a function that adds the first x powers of 1/2: f(x) = (1/2)^0 = 1 f(1) = (1/2)^0 + (1/2)^1 = 1+1/2 =3/2 f(2) = (1/2)^0 +(1/2)^1 + (1/2)^2 = 1+ 1/2+1/4=7/4

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

find a formula for f(x)

OpenStudy (amistre64):

a sequence of numbers is generated by the partial sums .... you are in effect making a new sequence that will hopefully be able to have an explicit formula. what are the next say, 3 partial sums?

OpenStudy (anonymous):

11/4 ?

OpenStudy (amistre64):

\[\frac{1}{1},\frac32,\frac{7}{4},\frac{15}{8}\]

OpenStudy (anonymous):

oh yeah.

OpenStudy (amistre64):

the next is double and add 1 ... 31/16 the next is double and add 1 63/64 .... seems to be the patter

OpenStudy (anonymous):

so how should I express that in terms of f(x) ?

OpenStudy (amistre64):

well, a few ways, so a little more context is in order. does it need to be a polynomial? can it be a summation rule? how do we need to define the rule?

OpenStudy (amistre64):

\[f_{n}=f_{n-1}+\frac{1}{2^n}\]

OpenStudy (amistre64):

we then determine the homogenous part, then work out the particular part

OpenStudy (amistre64):

does that make sense? or is that beyond the scope of your material?

OpenStudy (anonymous):

yes it is but your answer kind of makes sense, is that fx ? can't really see your subscript

OpenStudy (amistre64):

its fn at the moment ...

OpenStudy (anonymous):

Oh can it be a simple f(x) = and with a formula ? lol. i waasn't sure because it goes up to infinity.

OpenStudy (amistre64):

it can be, yes.

OpenStudy (anonymous):

hmm i think i figured it out, is it f(x) = 2 - (1/2)^x ? is that right ?

OpenStudy (amistre64):

that actually does work :) http://www.wolframalpha.com/input/?i=table+%5B2+-+%281%2F2%29%5En%5D%2C+n%3D0..8

OpenStudy (anonymous):

thanks ;D

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

okay so theres another part to this question lol

OpenStudy (anonymous):

use a graphing calculator to graph f(x) then make a conjecture about f(x) when x --> infinity

OpenStudy (amistre64):

you got a graping calc?

OpenStudy (anonymous):

yes, it goes in a straight line at y=3

OpenStudy (anonymous):

right ?

OpenStudy (amistre64):

2 - anything ... is not equal to 3. we would say that 3 is an upper bound, but it is not the smallest upper bound. for that matter, 12 is an upper bound :)

OpenStudy (amistre64):

(1/2)^x approaches zero as x approaches infinity .. 2-0 = ??

OpenStudy (anonymous):

2 ?

OpenStudy (amistre64):

yes, so if we graph it ... the line it approaches will be 2. we can conjecture that the limit of adding up this setup, for infinity ... will never exceed the value of 2. that is the smallest upper bound we can place on it.